OpenStudy (anonymous):

triangle QRS=triangle GHJ, RS=12, QR=10, QS=6, and HJ=2X-4. find x. a. 6 c. 7 b. 8 b. 10

OpenStudy (anonymous):

As we take QRS to GHJ, QR=GH RS=HJ QS=GJ (Take the = sign to be congruent signs) RS=12 HJ = RS = 12 = 2x-4 2x=12+4 2x=16 x=8

OpenStudy (anonymous):